Circular No. 5635

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1992 No. 3 G. S. Stringfellow and M. D. Gregg, Mount Stromlo and Siding Spring

Observatories, communicate: “We confirm this object to be a nova. Lowresolution spectra (range 490-940 nm) were obtained with the Boller & Chivens spectrograph on the 1.9-m telescope at Mt. Stromlo on Oct. 14.4 UT. Inspection of the raw spectra reveals strong Hα and O I (777.4-nm blend) emission lines. The Hα profile is quite asymmetric, having a sharp blue edge with no P-Cyg absorption and an extended red wing. The FWHM and FWZI of the Hα line yield 1200 and 3850 km/s, respectively. Moderate emission lines of O I (788.6, 823.5, 844.6 nm) and Na I (589.2 nm) all exhibit strong P-Cyg absorption (except for the weak O I 788.6-nm line), as does the O I 777.4-nm blend. Weak to moderate emission lines of Fe II multiplets (42, 49, 74, and 73) are present, indicating that this nova is a member of the ‘Fe II’ class. The nova appeared somewhat brighter on the TV monitor than was reported 24 hr earlier (IAUC 5634) and could be still rising or near maximum.”

R. H. McNaught, Anglo-Australian Observatory, reports his measurement from CCD frames taken Oct. 13 at the 1.0-m Siding Spring telescope: α = 18^{h}20^{m}39^{s}.43, δ = −23^{o}01′05′′.5 (equinox 1950.0).

Visual and photovisual magnitude estimates: Oct. 14.41 UT, 8.5 (P. Camilleri, Cobram, Victoria, Australia; T-Max 400 film); 14.71, 8.3 (D. Overbeek, Edenvale, R.S.A.); 14.77, 8.0 (H. Mikuz, Ljubljana, Slovenia).

NOVA CYGNI 1992 C. E. Woodward, University of Wyoming; and R. D. Gehrz and G. F.

Lawrence, University of Minnesota, report the following infrared magnitudes, obtained with the 2.3-m Wyoming Infrared Observatory telescope (+ bolometer detector) on Oct. 10.21 UT: J = 8.82 ± 0.10, H = 8.87 ± 0.08, K = 8.04 ± 0.12, L = 6.72 ± 0.05, M = 5.92 ± 0.05, N = 4.15 ± 0.09, [7.8 μm] = 2.70 ± 0.10, [8.7 μm] = 4.95 ± 0.10, [9.8 μm] = 5.30 ± 0.36 [10.3 μm] = 4.89 ± 0.34, [11.6 μm] = 4.61 ± 0.41, [12.5 μm] = 3.48 ± 0.17. Both the 12.5- and 7.8-μm magnitudes are well in excess of the continuum, suggesting that enhanced line emission from neon (IAUC 5497) is present. Further spectroscopic observations are encouraged.

1992 October 15 Daniel W. E. Green